# Measurement-aligned Flow for Inverse Problem

### v1

@@ 2.1 Conditional Diffusion and flow models
$dx=\left[(\dot{\alpha}_{t}-\alpha_{t}\xi_{t})\tilde{x}_{0}(t,x,y)+\xi_{t}x_{t}\right]dt+\epsilon_{t}dW_{t},$ (2)
where $\tilde{x}_{0}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]$ and $\xi_{t}=\frac{\dot{\sigma}_{t}}{\sigma_{t}}-\frac{\epsilon_{t}^{2}}{2\sigma_{t}^{2}}$. Then $x_{0}$ could be sampled by solving the SDE described in Eq. 2. Although different sampling methods can be chosen, generally, the iteration follows the form:
$x_{t-\Delta t}\sim\mathcal{N}(a_{t}\tilde{x}_{0}(t,x,y)+b_{t}x_{t},c_{t}^{2}\mathbb{I}).$ (3)
2.2 Training-free conditional inference
Given pretrained unconditional diffusion models $\hat{x}_{0}^{\theta}\approx\mathbb{E}[x_{0}\mid x_{t}]$, training-free approaches for conditional inference allows us to approximate the ideal denoiser $\tilde{x}_{0}(t,x_{t},y)$ (pokle2023training, ):
$\tilde{x}_{0}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]\approx\hat{x}_{0}^{\theta}+\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}\log p(y\mid x_{t}).$ (4)
For inverse problems or conditioning problems where $p(y\mid x_{0})$ is known but $p(y\mid x_{t})$ is intractable (chung2022diffusion, ; wu2024practical, ), denoising predictors $\hat{x}_{0}^{\theta}(t,x_{t})$ provide a convenient single-sample approximation for $p(y|x_{t})$, $p(y\mid x_{t})\approx p(y\mid\hat{x}_{0}^{\theta}(t,x_{t}))$. For example, $\Pi$GDM (song2023pseudoinverse, ) assumes $p_{t}(x_{0}\mid x_{t})\approx\mathcal{N}(\hat{x}_{0}^{\theta}(t,x_{t}),r_{t}^{2}\mathbb{I})$, then
[…]
Then the gradient guidance can be approximated by differentiating through the denoising predictor network using the chain rule:
$\nabla_{x_{t}}\log p(y\mid x_{t})\approx\left[(y-\hat{x}_{0}^{\theta})^{T}(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I})H\frac{\partial\hat{x}_{0}^{\theta}(t,x_{t})}{\partial x_{t}}\right]^{\mathsf{T}}.$ (6)
Substituting Eq. 6 into Eq. 4, we obtain an estimate of the ideal denoiser $\tilde{x}_{0}(t,x_{t},y)$. While this method works well in practice to produce the desired guidance effect (Sec. 5), it still relies on a crude approximation of the expectation in Eq. 5 and requires backpropagation through the denoising network. In this work, we present an alternative approach to estimate $\tilde{x}_{0}(t,x_{t},y)$ based on unconditional denoising predictors $\hat{x}_{0}^{\theta}(t,x_{t})$ and the measurement $y$.
3 Methodology
3.1 Measurement-Aligned Sampling for linear inverse problem
Measurement-Aligned Sampling aims to estimate the desired conditional denoiser $\tilde{x}_{0}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]$ at time $t$ by utilizing two measurements: an approximated unconditional denoiser $\hat{x}^{\theta}_{0}(t,x_{t})\approx\mathbb{E}[x_{0}\mid x_{t}]$ and the measurement $y$. MAS performs posterior sampling $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$ based on the following probabilistic model:
Prior: $\quad p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})\coloneqq\mathcal{N}(\hat{x}_{0}^{\theta},\rho_{t}^{2}\mathbb{I}_{d}),$ (7a)
[…]
where $\rho_{t}$ is a hyperparameter defining the prior variance.
Typically, computing the exact likelihood in Eq. 7b is infeasible since $p(x_{0}\mid\tilde{x}_{0})$ is unknown. $\Pi$GDM (song2023pseudoinverse, ) assume $p(x_{0}\mid\hat{x}_{0}^{\theta})=\mathcal{N}(\hat{x}_{0}^{\theta},r_{t}^{2}\mathbb{I})$, where $\hat{x}_{0}^{\theta}\approx\mathbb{E}[x_{0}\mid x_{t}]$. In our work, we revise this assumption to $p(x_{0}\mid\tilde{x}_{0})=\mathcal{N}(\tilde{x}_{0},r_{t}^{2}\mathbb{I})$. Consequently, for linear inverse problem where $p(y\mid x_{0})\coloneqq\mathcal{N}(Hx_{0},\sigma_{y}^{2}\mathbb{I})$, the likelihood $p(y\mid\tilde{x}_{0})$ has similar Gaussian form as in Eq. 5:
$p(y\mid\tilde{x}_{0})=\mathcal{N}(y;H\tilde{x}_{0},\sigma_{y}^{2}\mathbb{I}_{m}+r_{t}^{2}HH^{\mathsf{T}}).$ (8)
Given the prior $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})$ in Eq. 7a and the likelihood in Eq. 7b, the posterior distribution $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$ can be analytically derived via Bayesian linear regression, as detailed in Prop. 3.1. See App. A for the full derivation and proofs.
Proposition 3.1 (Bayesian Linear Regression).
If $p(y\mid\tilde{x}_{0})$ and $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})$ are both Gaussian or can be approximated as such, $p({y}\mid\tilde{x}_{0})=\mathcal{N}(H\tilde{x}_{0},\sigma_{y}^{2}{\mathbb{I}}+r_{t}^{2}HH^{\mathsf{T}})$, $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})=\mathcal{N}(\hat{x}_{0}^{\theta},\rho_{t}^{2}{\mathbb{I}_{d}})$, then $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)=\mathcal{N}(\mu_{\mathrm{post}},\Sigma_{\mathrm{post}})$:
$\Sigma_{\mathrm{post}}$ $=\left(\frac{1}{\rho_{t}^{2}}\mathbb{I}_{d}+H^{\mathsf{T}}(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I}_{m})^{-1}H\right)^{-1},$ (9)
$\mu_{\mathrm{post}}$ $=\Sigma_{\mathrm{post}}\left[\frac{1}{\rho_{t}^{2}}\hat{x}_{0}^{\theta}+H^{\mathsf{T}}(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I}_{m})^{-1}y\right].$ (10)
This result provides a concrete way to sample $\tilde{x}_{0}$ from the posterior $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})$ without backpropagation through the denoising network, as is needed in $\Pi$GDM (song2023pseudoinverse, ). We discuss implementation of this scheme before generalizing to non-Gaussian noise models in Sec. 4.
3.2 Implementation
Posterior sampling.
Given the posterior distribution $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$ explicitly derived via Bayesian linear regression in Prop. 3.1, we can efficiently sample $\tilde{x}_{0}\sim p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$. Defining $\eta_{1}\coloneqq{r_{t}^{2}}/{\rho_{t}^{2}}$, $\eta_{2}\coloneqq\sigma_{y}^{2}/\rho_{t}^{2}$, posterior sampling $\tilde{x}_{0}^{\text{ps}}\sim\mathcal{N}(\mu_{\text{post}},\Sigma_{\text{post}})$ can be implemented by a single draw of $\epsilon_{p}\sim\mathcal{N}(0,\mathbb{I}_{d})$:
$\tilde{x}_{0}^{\text{ps}}$ $=Y^{-1}[\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y+\rho_{t}\epsilon_{p}],\quad\epsilon_{p}\sim\mathcal{N}(0,\mathbb{I}_{d})$ (11)
$\text{where}\quad W$ $\coloneqq\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m},\quad Y\coloneqq\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H,$
then $\tilde{x}_{0}^{\text{ps}}\sim p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$.
Posterior mean.
In the implementation, posterior sampling in Eq. 11 introduces additional noise $\rho_{t}Y^{-1}\epsilon_{p}$, which leads to noisy restoration in the experiment, See App. G for more details. For simplicity and improved reconstruction quality, we could directly replace $\tilde{x}_{0}^{\text{ps}}$ in Eq. 11 by the posterior mean $\tilde{x}_{0}^{\text{pm}}\coloneqq\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]$:
$\tilde{x}_{0}^{\text{pm}}=Y^{-1}[\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (12)
Efficient linear solves via the SVD of $H$.
Computing the inverse $W^{-1}$ and $Y^{-1}$ in Eq. 11 and Eq. 12 naively can be computationally expensive. Thus, in practice, we would like to avoid directly computing the inverse. Let $H=U\Sigma V^{\mathsf{T}}$ with singular values $s_{1},\dots,s_{n}$. Then
[…]
see App. A for the proof. Hence both $W^{-1}v$ and $Y^{-1}u$ reduce to inexpensive diagonal scalings in the SVD basis, avoiding the calculation of any explicit matrix inversion or square-root.
3.3 Exploring alignment trade-offs via an optimization perspective
We can interpret the posterior mean $\tilde{x}_{0}^{\text{pm}}$ from an optimization perspective to explicitly see how our approach incorporates alignment from the prior and the measurement. Specifically, $\tilde{x}_{0}^{\text{pm}}$ minimizes the following objective function (see App. A for the proof):
$\tilde{x}_{0}^{\text{pm}}=\arg\min\{\|\tilde{x}_{0}-\hat{x}_{0}^{\theta}\|^{2}+\|y-H\tilde{x}_{0}\|_{W^{-1}}^{2}\}$ (14)
where $W=\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m}$ (with $\eta_{1}\geq 0,\eta_{2}\geq 0$) emerges as a metric that reflects the relative importance of measurement fidelity and prior regularization. As $\eta_{1}=0,\eta_{2}>0$, Eq. 14 reduced to:
$\tilde{x}_{0}^{\text{pm}}=\arg\min\{\|\tilde{x}_{0}-\hat{x}_{0}^{\theta}\|^{2}+\frac{1}{\eta_{2}}\|y-H\tilde{x}_{0}\|^{2}\}$ (15)
This reduced form represents a standard Tikhonov regularization (ridge regression) problem. The choice of $\eta_{2}$ here controls the trade-off between fidelity to the measurements $y$ and $\hat{x}_{0}^{\theta}$. While as $\eta_{1}>0,\eta_{2}=0$, the optimization takes on the following generalized least squares (GLS) form:
$\tilde{x}_{0}^{\text{pm}}=\arg\min\{\|\tilde{x}_{0}-\hat{x}_{0}^{\theta}\|^{2}+\frac{1}{\eta_{1}}\|y-H\tilde{x}_{0}\|_{(HH^{\mathsf{T}})^{-1}}^{2}\}$ (16)
Here, the weighting term $(HH^{\mathsf{T}})^{-1}$ indicates a Mahalanobis-type distance metric, emphasizing directions corresponding to smaller singular values of $H$. The choice of $\eta_{1}$ and $\eta_{2}$ significantly affects the reconstruction quality, as will be demonstrated experimentally.
Remark 1 (Connection with DAPS (zhang2024improving, )).
As $\eta_{1}=0$, $\tilde{x}_{0}^{\text{ps}}$ is reduced to posterior sampling in DAPS.
Remark 2 (Connection with DDNM (wang2022zero, )).
As $\eta_{2}=0$ and $\eta_{1}\rightarrow 0$, $\tilde{x}_{0}^{\text{pm}}\rightarrow\tilde{x}_{0}^{\text{DDNM}}=\hat{x}_{0}^{\theta}+H^{\dagger}(y-H\hat{x}_{0}^{\theta})$. Thus, in this limiting case, MAS recovers DDNM.
Figure 2: 2D illustration of the influence of parameters $\eta_{1}$ and $\eta_{2}$. Dots represent posterior means $\tilde{x}_{0}^{\text{pm}}$, calculated via Eq. 12. (a) Parameter $\eta_{1}$ controls the trade-off between $\hat{x}_{0}^{\theta}$ and $\tilde{x}_{0}^{\text{DDNM}}$: as $\eta_{1}\rightarrow\infty$, the posterior mean $\tilde{x}_{0}^{\text{pm}}$ approaches $\hat{x}_{0}^{\theta}$; as $\eta_{1}\rightarrow 0$, it converges to $\tilde{x}_{0}^{\text{DDNM}}$. (b) Adjusting $\eta_{2}$ alters the posterior trajectory differently from varying $\eta_{1}$.
Remark 3 (Overshooting trick).
Theoretically, $\eta_{1}\geq 0$ since $\eta_{1}\coloneqq r_{t}^{2}/\rho_{t}^{2}$, however, the posterior mean in Eq. 12 allows negative $\eta_{1}$. As shown in Fig. 2, negative $\eta_{1}$ leads to overshooting, which demonstrates better performance than DDNM in the experiments.
4 Generalizing to noisy inverse problems
4.1 Identifying the introduced noise
To illustrate our ideas, we first consider a simple forward model: $y=x_{0}+\epsilon_{y}$. Calculating the posterior mean $\tilde{x}_{0}^{\text{pm}}$ (Eq. 12) leads to:
$\tilde{x}_{0}^{\text{pm}}=\hat{x}_{0}^{\theta}+\frac{y-\hat{x}_{0}^{\theta}}{\eta_{1}+\eta_{2}+1}=\hat{x}_{0}^{\theta}+\frac{x_{0}-\hat{x}_{0}^{\theta}}{\eta_{1}+\eta_{2}+1}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\epsilon_{y}}{\eta_{1}+\eta_{2}+1}},$ (17)
where $\epsilon_{y}/(\eta_{1}+\eta_{2}+1)$ is the additional noise introduced to $\tilde{x}_{0}^{\text{pm}}$, and will be further introduced into $x_{t-\Delta t}$. To address this issue, we propose two approaches for addressing known Gaussian noise (Sec. 4.2) and unknown noise Sec. 4.3.
4.2 Addressing Gaussian noise with known variance
To handle Gaussian noise with known variance, we modify Eq. 17 and Eq. 3 as:
$\tilde{x}_{0}=\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{t}}\frac{y-\hat{x}_{0}^{\theta}}{\eta_{1}+\eta_{2}+1},\quad x_{t-\Delta t}\sim\mathcal{N}(a_{t}\tilde{x}_{0}+b_{t}x_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t}}\mathbb{I}).$ (18)
Here $\lambda_{t}$ and $r_{t}$ are two parameters that can control the total noise introduced to $x_{t-\Delta t}$. In our work, we adopt similar two principles as DDNM$+$ (wang2022zero, ) for handling Gaussian noise: (i) the total noise introduced in $x_{t-\Delta t}$ should be $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$ to conform the correct distribution of $x_{t-\Delta t}$ in Eq. 3; (ii) $\lambda_{t}$ should be as close to $1$ as possible to maximize the preservation of $\tilde{x}_{0}^{\text{pm}}$. As $\epsilon_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$, principle (i) and principle (ii) are equivalent to:
$\left(\frac{a_{t}\lambda_{t}\sigma_{y}}{\eta_{1}+\eta_{2}+1}\right)^{2}+\gamma_{t}$ $=c_{t}^{2},\quad\lambda_{t}=\begin{cases}1,&c_{t}\geq\dfrac{a_{t}\sigma_{y}}{\eta_{1}+\eta_{2}+1}\\[4.0pt] \dfrac{c_{t}(\eta_{1}+\eta_{2}+1)}{a_{t}\sigma_{y}},&c_{t}<\dfrac{a_{t}\sigma_{y}}{\eta_{1}+\eta_{2}+1}\end{cases}.$ (19)
The more accurate version for general form of $H$ could be found in App. B. Note that the revision didn’t introduce additional parameters.
@@ 4.3 Addressing unknown noise and non-differentiable measurements
Figure 3: The sample process of solving inverse problem with unknown noise. Here we set $\eta_{1}=0$ and $\eta_{2}=0.5a_{t}/c_{t}$.
Addressing unknown noise or non-Gaussian noise. When the measurement noise $\sigma_{y}$ is non-Gaussian or unknown, it becomes difficult to ensure that the total noise in $x_{t-\Delta t}$ follows the desired distribution $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$, To address this, we continue to sample $x_{t-\Delta t}$ using Eq. 3. Next, the noise introduced to $x_{t-\Delta t}$ is the sum of two components:
@@
Here $\epsilon_{\text{ng}}$ is related to the noise introduced by unknown noise $\epsilon_{y}$, while $\epsilon_{\text{g}}$ is the noise added by the diffusion process. To minimize the effect of unknown noise $\epsilon_{\text{ng}}$, it is desirable for $\eta_{1}+\eta_{2}+1$ to be sufficiently large. However, smaller values of $\eta_{1}$ and $\eta_{2}$ result in better consistency with the measurement $y$, as illustrated in Fig. 2. To balance this trade-off, we propose using a small $\eta_{1}+\eta_{2}$ during the early stages of sampling to fully exploit measurement information. As sampling progresses, $\eta_{1}+\eta_{2}$ should be gradually increased to suppress the impact of $\epsilon_{\text{ng}}$. The underlying intuition is that, in the early sampling stage, $x_{t}$ is still highly noisy and $a_{t}\approx 0$, so the influence of $\epsilon_{\text{ng}}$ is negligible even when $\eta_{1}+\eta_{2}$ is small. As shown in Fig. 3, $\tilde{x}^{\text{pm}}$ is initially more aligned with the degraded observation, but progressively shifts toward $\hat{x}_{0}^{\theta}$ as sampling evolves.
For a general degradation operator $H$, we recommend setting $\eta_{2}=ka_{t}/c_{t}$, where $k$ is a constant determined by the characteristics of the introduced noise. The rationale behind this design choice is further detailed in App. B.
Addressing non-differentiable measurements. For solving inverse problem with non-differentiable measurements such as JPEG restoration and quantization, the degraded images can be viewed as "noisy images" with unknown noise, modeled by $y=x+\epsilon_{y}$. In these scenarios, our proposed strategy naturally extends by treating the unknown degradations as implicit noise.
@@ 5.1 Experimental setup
Pretrained models and baselines. We utilize the pre-trained checkpoint (chung2022diffusion, ) on the FFHQ dataset and the pre-trained checkpoint (dhariwal2021diffusion, ) on the Imagenet dataset. We compare our methods with the following baselines: DDNM (wang2022zero, ), DDRM (kawar2022denoising, ), $\Pi$GDM (song2023pseudoinverse, ), RedDiff (mardani2023variational, ), DAPS (zhang2024improving, ).
@@
Tasks. (1) We evaluate performance on the following linear inverse problems: super-resolution (pooling filter), deblurring (uniform kernel of size $25$), inpainting (box mask), inpainting (freeform mask), and colorization. (2) We consider two unknown noise types: salt-and-pepper noise ($10\%$ pixels set randomly to ±$\pm 1$) and periodic noise (sinusoidal pattern with amplitude $0.2$ and frequency $5$). (3) We address JPEG restoration with quality factors $\mathrm{QF}=2$ and $\mathrm{QF}=5$. (4). For quantization, we consider the challenging case of 2-bit quantization.
@@ 5.2 Ablation study
Figure 4: Ablation study of $\eta_{1}$ and $\eta_{2}$ on solving super-resolution and deblurring.
@@
We conduct ablation studies on parameters $\eta_{1}$ and $\eta_{2}$ using two inverse problems: super-resolution (noise-free, $\epsilon_{y}=0$) and deblurring (noisy, $\epsilon_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$). Results presented in Fig. 4 demonstrate that for noise-free super-resolution, the highest PSNR and lowest FID scores are achieved by setting $\eta_{2}=0$ and a negative $\eta_{1}=-0.4$. This indicates that appropriate "overshooting" enhances restoration quality. For the noisy deblurring task, negative $\eta_{2}$ similarly improves PSNR and lowers the FID score, further confirming the benefit of overshooting. Nevertheless, optimal performance is observed at $\eta_{2}=0.01$, suggesting that suitable regularization via $\eta_{2}$ significantly enhances reconstruction results. Additionally, we observe a performance drop as $\eta_{2}=0$ and $\eta_{1}\rightarrow 0$ (where $\tilde{x}_{0}^{\mathrm{bm}}\approx\tilde{x}_{0}^{\mathrm{DDNM}}$).
@@
Table 2: Quantitative evaluation of solving noisy linear inverse problems with Gaussian noise (known variance, $\sigma_{y}=0.2$).
Task Method NFE FFHQ ImageNet
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
SR 8×8$\times$8 × DDNM 20 29.87 0.258 0.837 108.24 27.71 0.385 0.711 189.57
DDRM 20 29.53 0.269 0.834 114.73 27.40 0.397 0.704 192.43
$\Pi$GDM 100 21.21 0.632 0.384 369.61 19.66 0.604 0.284 303.60
RED-Diff 100 24.06 0.604 0.604 320.15 23.22 0.61 0.522 310.29
DAPS 100 18.71 0.622 0.370 328.65 17.18 0.658 0.275 309.68
MAS (ours) 20 29.88 0.254 0.838 107.67 27.71 0.377 0.712 185.43
Deblurring DDNM 20 21.31 0.571 0.173 175.71 19.98 0.592 0.134 295.76
DDRM 20 29.25 0.271 0.823 121.00 27.54 0.405 0.704 213.13
$\Pi$GDM 100 5.15 0.846 0.004 493.09 - - - -
RED-Diff 100 5.20 0.845 0.004 502.66 5.20 0.804 0.004 457.13
DAPS 100 19.01 0.616 0.365 244.10 16.60 0.668 0.231 307.79
MAS (ours) 20 29.62 0.267 0.827 119.43 27.88 0.408 0.710 218.51
Inpaint (Freeform) DDNM 20 25.99 0.191 0.887 117.12 27.04 0.218 0.854 98.60
DDRM 20 26.53 0.181 0.903 93.73 27.28 0.205 0.855 75.71
$\Pi$GDM 100 24.88 0.396 0.504 129.45 25.01 0.352 0.513 119.64
RED-Diff 100 25.22 0.404 0.517 174.56 25.47 0.373 0.538 150.64
DAPS 100 18.99 0.471 0.330 148.37 19.43 0.419 0.385 138.87
MAS (ours) 20 27.20 0.157 0.913 92.56 27.92 0.177 0.871 77.61
Inpaint (Box) DDNM 20 27.87 0.165 0.880 94.27 24.88 0.247 0.826 168.99
DDRM 20 28.36 0.159 0.896 84.76 24.87 0.236 0.827 151.55
$\Pi$GDM 100 25.72 0.375 0.509 153.81 23.44 0.377 0.486 151.99
RED-Diff 100 25.87 0.382 0.522 186.75 24.19 0.409 0.518 239.49
DAPS 100 19.86 0.437 0.338 163.03 18.03 0.448 0.354 212.95
MAS (ours) 20 28.77 0.135 0.904 83.79 24.75 0.213 0.838 143.01
Colorization DDNM 20 28.41 0.216 0.937 113.04 27.08 0.315 0.877 162.19
DDRM 20 28.34 0.220 0.930 108.50 27.19 0.285 0.876 124.94
$\Pi$GDM 100 23.51 0.552 0.409 185.22 22.57 0.513 0.453 216.30
RED-Diff 100 24.50 0.557 0.403 206.85 23.67 0.537 0.462 274.81
DAPS 100 21.71 0.491 0.479 145.99 20.45 0.492 0.531 185.29
MAS (ours) 20 28.43 0.205 0.941 112.94 27.39 0.282 0.893 147.07

### v2

@@ 4.2 Addressing Gaussian noise with known variance
Derivations for more general forms of $H$ can be found in App.˜B. Note that the revision does not introduce additional parameters.
@@ 4.3 Addressing unknown noise and non-differentiable measurements
Figure 3: The sample process of solving inverse problems with unknown noise, where $\hat{x}_{0}^{\theta}\approx m_{0\mid t}$ is the denoising output. Here we set $\eta_{1}=0$ and $\eta_{2}=0.5a_{t}/c_{t}$.
Addressing unknown noise or non-Gaussian noise. When the measurement noise $\sigma_{y}$ is non-Gaussian or unknown, it becomes difficult to ensure that the total noise in $x_{t-\Delta t}$ follows the desired distribution $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$, To address this, we continue to sample $x_{t-\Delta t}$ using Eq.˜2. Next, the noise introduced to $x_{t-\Delta t}$ is the sum of two components:
@@
Here $\epsilon_{\text{ng}}$ is related to the noise introduced by unknown noise $\epsilon_{y}$, while $\epsilon_{\text{g}}$ is the noise added by the diffusion process. To minimize the effect of unknown noise $\epsilon_{\text{ng}}$, it is desirable for $\eta_{1}+\eta_{2}+1$ to be sufficiently large. However, smaller values of $\eta_{1}$ and $\eta_{2}$ result in better consistency with the measurement $y$, as illustrated in Fig.˜2. To balance this trade-off, we propose using a small $\eta_{1}+\eta_{2}$ during the early stages of sampling to fully exploit measurement information. As sampling progresses, $\eta_{1}+\eta_{2}$ should be gradually increased to suppress the impact of $\epsilon_{\text{ng}}$. The underlying intuition is that, in the early sampling stage, $x_{t}$ is still highly noisy and $a_{t}\approx 0$, so the influence of $\epsilon_{\text{ng}}$ is negligible even when $\eta_{1}+\eta_{2}$ is small. As shown in Fig.˜3, $x_{0}^{*}$ is initially more aligned with the degraded observation, but progressively shifts toward $m_{0\mid t}$ as sampling evolves.
For a general degradation operator $H$, we recommend setting $\eta_{2}=ka_{t}/c_{t}$, where $k$ is a constant determined by the characteristics of the introduced noise. The rationale behind this design choice is further detailed in App.˜B.
Addressing non-differentiable measurements. For solving inverse problems with non-differentiable measurements such as JPEG restoration and quantization, the degraded images can be viewed as "noisy images" with unknown noise, modeled by $y=x+\epsilon_{y}$. In these scenarios, our proposed strategy naturally extends by treating the unknown degradations as implicit noise.
@@ 5.1 Experimental setup
[…]
@@
Tasks. (1) We evaluate performance on the following linear inverse problems: super-resolution (bicubic filter), deblurring (uniform kernel of size $9$), inpainting (with a box mask), inpainting (with a 70% random mask), and colorization. (2) We consider two unknown noise types: salt-and-pepper noise ($10\%$ pixels set randomly to $\pm 1$) and periodic noise (sinusoidal pattern with amplitude $0.2$ and frequency $5$). (3) We address JPEG restoration with quality factors $\mathrm{QF}=2$ and $\mathrm{QF}=5$. (4). For quantization, we consider the challenging case of 2-bit quantization.
@@ 5.2 Ablation study
Figure 4: Ablation study of $\eta_{1}$ and $\eta_{2}$ on solving super-resolution and deblurring. We set NFE=20 for all tasks.
@@
We conduct ablation studies on parameters $\eta_{1}$ and $\eta_{2}$ using two inverse problems: super-resolution (noise-free, $\epsilon_{y}=0$) and deblurring (noisy, $\epsilon_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$). Results presented in Fig.˜4 demonstrate that for noise-free super-resolution, the highest PSNR and lowest FID scores are achieved by setting $\eta_{2}=0$ and a negative $\eta_{1}=-0.45$. This indicates that appropriate "overshooting" enhances restoration quality. For the noisy deblurring task, negative $\eta_{2}$ yields an improvement of more than 0.5 in PSNR and a reduction of over 5 in FID, further confirming the benefit of overshooting.
@@
Table 2: Quantitative evaluation of solving image restoration FFHQ (left) and ImageNet (right), with Gaussian noise (known variance, $\sigma_{y}=0.05$).
Task Method FFHQ ImageNet
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
SR $4\times$ DPS $25.86$ $0.753$ $0.269$ $81.07$ $21.13$ $0.489$ $0.361$ $106.32$
DDRM $26.58$ $0.782$ $0.282$ $79.25$ $22.62$ $0.521$ $0.324$ $103.85$
DDNM $28.03$ $0.795$ $0.197$ $64.62$ $23.96$ $0.604$ $0.475$ $98.62$
DCDP $28.66$ $0.807$ $0.178$ $53.81$
FPS-SMC $28.42$ $0.813$ $0.204$ $49.25$ $24.82$ $0.703$ $0.313$ $97.51$
DiffPIR $26.64$ $0.260$ $65.77$ $23.18$ $0.371$ $106.32$
RED-Diff $28.63$ $0.748$ $0.288$ $126.78$ $25.43$ $0.639$ $0.336$ $153.37$
DAPS $29.07$ $0.818$ $0.177$ $51.44$ $25.89$ $0.694$ $0.276$ $83.57$
MAS $\mathbf{30.56}$ $\mathbf{0.865}$ $\mathbf{0.131}$ $61.38$ $\mathbf{27.20}$ $\mathbf{0.751}$ $\mathbf{0.215}$ $88.61$
Inpaint (Box) DPS $22.51$ $0.792$ $0.209$ $61.27$ $18.94$ $0.722$ $0.257$ $126.52$
DDRM $22.26$ $0.801$ $0.207$ $78.62$ $18.63$ $0.733$ $0.254$ $116.37$
DDNM $24.47$ $0.837$ $0.235$ $46.59$ $21.64$ $0.748$ $0.319$ $103.97$
DCDP $23.89$ $0.760$ $0.163$ $45.23$
FPS-SMC $24.86$ $0.823$ $0.146$ $48.34$ 22.16{22.16} $0.726$ $0.208$ $111.58$
RED-Diff $24.68$ $0.767$ $0.175$ $86.78$ $21.32$ $0.728$ $0.247$ $123.55$
DAPS $24.07$ $0.814$ $0.133$ $43.10$ $21.43$ $0.725$ $0.214$ $109.85$
MAS $\mathbf{24.95}$ $\mathbf{0.879}$ $\mathbf{0.082}$ $\mathbf{37.67}$ $21.15$ $\mathbf{0.817}$ $\mathbf{0.168}$ $\mathbf{95.96}$
Inpaint (Random) DPS $25.46$ $0.823$ $0.203$ $69.20$ $23.52$ $0.745$ $0.297$ $87.53$
DDNM $29.91$ $0.817$ $0.121$ $44.37$ 31.16{31.16} $0.841$ $0.191$ $63.84$
DCDP $30.69$ $0.842$ $0.142$ $52.51$
FPS-SMC $28.21$ $0.823$ $0.261$ $61.23$ $24.52$ $0.701$ $0.316$ $79.12$
RED-Diff $29.73$ $0.814$ $0.200$ $104.19$ $27.04$ $0.753$ $0.226$ $92.24$
DAPS $31.12$ $0.844$ $0.098$ $32.17$ $28.44$ $0.775$ $0.135$ $54.25$
MAS $\mathbf{33.10}$ $\mathbf{0.923}$ $\mathbf{0.073}$ 34.75{34.75} $\mathbf{29.05}$ $\mathbf{0.838}$ $\mathbf{0.113}$ $\mathbf{30.19}$
Deblurring (Uniform) DDNM $26.58$ $0.704$ $0.210$ $68.83$ $25.69$ $0.630$ $0.261$ $83.63$
DDRM $29.19$ $0.835$ $0.172$ $87.12$ $26.31$ $0.711$ $0.267$ $118.36$
DAPS $28.92$ $0.758$ $0.204$ $76.57$ $25.43$ $0.616$ $0.293$ $103.55$
MAS $\mathbf{30.58}$ $\mathbf{0.857}$ $0.174$ $103.88$ $26.25$ $0.700$ $0.295$ $141.58$
Color DDNM $24.83$ $0.868$ $0.244$ $85.15$ $22.57$ $0.884$ $0.271$ $87.48$
DDRM $23.27$ $0.881$ $0.250$ $100.48$ $21.12$ $0.819$ $0.346$ $103.39$
RED-Diff $24.21$ $0.785$ $0.304$ $107.64$ $22.18$ $0.782$ $0.368$ $104.40$
DAPS $23.92$ $0.825$ $0.263$ $88.09$ $22.13$ $0.830$ $0.323$ $89.30$
MAS $24.23$ $\mathbf{0.919}$ $\mathbf{0.187}$ $\mathbf{72.33}$ $\mathbf{22.66}$ $\mathbf{0.886}$ $\mathbf{0.258}$ $\mathbf{83.17}$